# Fast3R: Towards 3D Reconstruction of 1000+ Images in One Forward Pass

### v1

@@ Appendix D Bundle Adjustment (via Gaussian splatting)
While not necessary, using bundle-adjustment at inference time can also improve Fast3R’s performance. We show an example of bundle adjustment using Gaussian Splatting (GS-BA).
Specifically, we use InstantSplat [17] to optimize a set of gaussians per-scene, using initializations from a pointcloud, and updates the locations and poses in order to minimize reprojection error. We show an example of the Gaussian reconstruction in Figure 13 shows an example reconstruction on CO3D.
@@
We can compare against ground-truth trajectories from COLMAP. We found that GS-BA significantly reduces both the pose and translation error. Table 6 quantifies this, showing over a 2.5x reduction in translation error and 4x redution in rotational error on the “Family” scene from Tanks and Temples, which we found to be representative. We show a visualization of the original reconstruction and the poses pre- and post- bundle adjustment. There are only 8 scenes in the evaluation set in InstantSplat.
@@
Appendix E More Visualizations
@@
In Figure 15, we visualize the point cloud produced using local vs. global point maps.
Figure 15: Effect of using local vs. global point map. Global point maps provide good anchors for locations of points while local point maps use those anchors (by aligning using ICP on the anchor points to the global point map) to provide more accurate point locations.

### v2

@@ Appendix D Bundle Adjustment (via Gaussian Splatting)
While not necessary, using bundle adjustment at inference time can also improve Fast3R’s performance. We show an example of bundle adjustment using Gaussian Splatting (GS-BA).
Specifically, we use InstantSplat [16] to optimize a set of Gaussians per scene, using initializations from a point cloud, and update the locations and poses in order to minimize reprojection error. We show an example of the Gaussian reconstruction in Figure 11 shows an example reconstruction on CO3D.
@@
We can compare against ground-truth trajectories from COLMAP. We found that GS-BA significantly reduces both the pose and translation error. Table 6 quantifies this, showing over a 2.5x reduction in translation error and a 4x reduction in rotational error on the “Family” scene from Tanks and Temples, which we found to be representative. We show a visualization of the original reconstruction and the poses pre- and post-bundle-adjustment. There are only 8 scenes in the evaluation set in InstantSplat.
@@
Methods ScanNet ETH3D DTU T&T
rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$ rel $\downarrow$ $\tau\uparrow$
COLMAP-DENSE 38.0 22.5 89.8 23.2 20.8 69.3 25.7 76.4
DUSt3R 224 5.86 50.84 4.71 61.74 2.76 77.32 5.54 56.38
DUSt3R 512 4.93 60.20 2.91 76.91 3.52 69.33 3.17 76.68
Fast3R 6.27 50.27 4.68 62.68 3.92 62.60 4.43 63.95
Table 7: Multi-view depth evaluation. DUSt3R and Fast3R perform on par, while significantly outperforming COLMAP-DENSE.
Appendix E Multi-view Depth Evaluation
We compare Fast3R (using the local pointmap prediction) with DUSt3R and COLMAP on multi-view depth estimation tasks and show results in Table 7.
Appendix F More Visualizations
@@
Figure 13: Large-scale reconstruction: Spann3R vs. Fast3R on the Lighthouse scene from Tanks & Temples dataset.
Figure 14: Effect of using local vs. global pointmap. Global point maps provide good anchors for locations of points while local point maps use those anchors (by aligning using ICP on the anchor points to the global point map) to provide more accurate point locations. Best viewed when zoomed in.
Figure 15: Visualizations of results on NRGBD scenes. Fast3R learns the regularity of indoor rooms (square-like shapes) and demonstrates loop closure capabilities.
F.1 4D Reconstruction: Qualitative Results